OpenStudy (anonymous):

\[4y''-4y'+y=0\] y(0)=1 y'(0)=-1.5 auxiliary equation \[4r^2-4r+=0\] \[(2r-1)(2r-1)=0\] \[r_1=.5\] \[r_2=.5\] \[y(x)=c_1e^{.5x}+c_2e^{.5x}\] \[y(0)=c_1e^{.5x}+c_2e^{.5x}=0\] \[y'(0)=c_10.5e^{0.5x}+c_20.5e^{0.5x}=-1.5\]

OpenStudy (turingtest):

so you are having trouble finding c1 and c2?

OpenStudy (lgbasallote):

wont it be better to use 1/2 instead of 0.5?

OpenStudy (anonymous):

yeah :S

OpenStudy (lgbasallote):

..just saying

OpenStudy (turingtest):

you did not plug in x=0 in you particulars

OpenStudy (anonymous):

forgot to do that

OpenStudy (lgbasallote):

you also equated y(0) to 0

OpenStudy (lgbasallote):

should be y(0) = 1

OpenStudy (anonymous):

\[y(0)=c_1e^{.5(0)}+c_2e^{.5(0)}=1\] \[y(0)=c_1+c_2=1\]

OpenStudy (turingtest):

\[y(0)=c_1e^{0}+c_2e^{0}=c_1+c_2=1\]\[y'(0)=0.5c_10+0.5c_2=-1.5\]

OpenStudy (anonymous):

repeated roots mean the solution looks like C1*e^lambda*x + C2*x*e^lambda*x

OpenStudy (turingtest):

good call, I missed that

OpenStudy (turingtest):

so\[y(0)=c_1e^{0}+c_2(0)e^{0}=c_1=1\]\[y'(0)=0.5(1)+0.5c_2=-1.5\]now it's easy

OpenStudy (anonymous):

why is \[c_2(0)e^0\] why do you have it multiplied by zero

OpenStudy (turingtest):

because of the repeated root you should have had\[y(x)=c_1e^{.5x}+c_2xe^{.5x}\]

OpenStudy (turingtest):

from which the above follows

OpenStudy (anonymous):

oh I see.

OpenStudy (anonymous):

and now it's just a matter of substitution Thanks!!!!

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

No problem.